OpenStudy (anonymous):

√(x-95)+9=6 solve for x and identify if it is an extraneous solution

OpenStudy (the_fizicx99):

Do you have an idea on how to start using my previous post as a guide?

OpenStudy (the_fizicx99):

@yamaliz95

OpenStudy (anonymous):

no i don't

OpenStudy (the_fizicx99):

Hm, k, start by subtracting 9 from both sides

OpenStudy (anonymous):

ok i know what the x value is. i just don't know how to find if it's extraneous or not

OpenStudy (the_fizicx99):

Ok extraneous basically means it doesn't work for the equation, like if you after plug in and simplify and it gives you something like 11 = 25 then it doesn't work. So what's the x value you have?

OpenStudy (anonymous):

the x value i have is 104

OpenStudy (the_fizicx99):

K, so plug in 104 for x, √(104-95)+9=6 √(9) + 9 = 6 What is the square root of 9? (simple) >.>

OpenStudy (anonymous):

i got that too. so now i know that it's extraneous thank you..

OpenStudy (the_fizicx99):

Yw, 3 + 9 = 6 12 = 6 therefore it doesn't work :D Gl!